OpenStudy (anonymous):

what is the value of b x^2+y^2+ax+by+c=0

OpenStudy (helder_edwin):

the value of \(b\) for what?

OpenStudy (anonymous):

for that equation, \[x ^{2}+y ^{2^{}}+ax+by+c=0\]

OpenStudy (anonymous):

@helder_edwin

OpenStudy (helder_edwin):

r u sure the question is exactly that? it kind of doesn't make sense

OpenStudy (anonymous):

this graph is part of the question

OpenStudy (helder_edwin):

u also need to find the values of a, b, and c so that the equation describes the circle of the figure. complete both squares. u know the center and the radius of the circle (see the graph)

OpenStudy (anonymous):

that doesnt help me

OpenStudy (anonymous):

but thanks

OpenStudy (helder_edwin):

do u know how to complete the square?

OpenStudy (anonymous):

no

OpenStudy (helder_edwin):

ok i will do it. first the center is C(-2,1) and r=3, do u agree?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I got that far.

OpenStudy (helder_edwin):

ok know, the equation of a circle should be \[\large (x-h)^2+(y-k)^2=r^2 \] we have to make the equation \[\large x^2+y^2ax+by+c=0 \] to resemble this. OK?

OpenStudy (helder_edwin):

\[\large x^2+y^2+ax+by+c=0 \]

OpenStudy (anonymous):

(x+2)^2+(y-1)^2=3^2

OpenStudy (helder_edwin):

we know that h=-2, k=1, and r=3, agree?

OpenStudy (helder_edwin):

yes. expand that!

OpenStudy (helder_edwin):

can u do it?

OpenStudy (anonymous):

(x+2)+(x+2)+(y-1)+(y-1)=9?

OpenStudy (anonymous):

I'm so confuseddd.

OpenStudy (helder_edwin):

no \[\large (x+2)^2+(y-1)^2=3^2 \] \[\large (x^2+4x+4)+(y^2-2x+1)=9 \] \[\large x^2+y^2+4x-2y+5=9 \] \[\large x^2+y^2+4x-2y-4=0 \]

OpenStudy (helder_edwin):

got it?

OpenStudy (anonymous):

so b is 4?

OpenStudy (helder_edwin):

no, b is the coefficient of \(y\)

OpenStudy (anonymous):

oh -2?

OpenStudy (helder_edwin):

yes

OpenStudy (anonymous):

can you help me with this one?

OpenStudy (helder_edwin):

same question?

OpenStudy (anonymous):

have to find a now. so the center is (-1,3) and radius is 4 right?

OpenStudy (anonymous):

(x+1)+(y-3)=16?

OpenStudy (raden):

correction number one, the centre of circle is (1,-2) not (-2,1). so, h = 1 and k = -2 now remember the equation of circle is x^2 + y^2 +ax + by + c = 0 or (x - h)^2 + (y -k)^2 = 0 h,k is the centre of circle. the relation of b and k is b = -1/2(k) so, b = -1/2(-2) = 1

OpenStudy (helder_edwin):

YES. I am so sorry, my mistake.

OpenStudy (anonymous):

oh boy. okay.

OpenStudy (helder_edwin):

in the second case the center is (3,-1) and r=4

OpenStudy (raden):

correction again, im not a boy, but a father of that boy ;p

OpenStudy (anonymous):

does A have a certain formula that can be used to find it?

OpenStudy (helder_edwin):

dont learn formulas, learn concepts. do what i did before: expand \[\large (x-3)^2+(y+1)^2=4^2 \] it should take a minute

OpenStudy (snuggielad):

Welcome To Openstudy

OpenStudy (raden):

a = -1/2 * h

OpenStudy (anonymous):

so (x^2-6x+9)+(y^2+2x+1)=16?

OpenStudy (anonymous):

so a would be = -3 1/2

OpenStudy (helder_edwin):

a=-6, that is what u got.

OpenStudy (ranga):

It should be (x^2-6x+9)+(y^2+2y+1)=16 (you had 2x, it should be 2y). In bx^2+y^2+ax+by+c=0, a is the coefficient of x. The coefficient of x in the equation you got is -6 So a = -6.